# Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

### v3

@@ Abstract
Large language models are good knowledge bases but struggle to perform equally well for all classes in simple text classification tasks. This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies. We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others, for both base LLMs and more sophisticatedly trained, larger chat LLMs. This class accuracy imbalance is difficult to solve from the root via “better” pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time optimization. To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update. Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve. We perform extensive evaluations on three LLMs across seven NLP classification tasks, in different prompting settings. Results show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
@@ 1 Introduction
The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience. Not mitigating such imbalances may lead to serious damages in some cases. For example, a patient can trust an LLM about drug-drug interaction type predictions without knowing that the model over-predicts “no interaction” and has low accuracies for other categories of interactions. If the patient trusts “no interaction” when the ground truth is an actual drug interaction type, taking the drugs together can put the patient’s life at risk. Therefore, it is high time we took actions to reduce class accuracy imbalances, better without hurting overall accuracy.
It is difficult to disentangle all the factors that cause class accuracy imbalance, including biases incurred in pre-training, supervised fine-tuning, RLHF, prompting specifics, etc. Therefore, it is challenging to comprehensively resolve the root causes through pre-training or fine-tuning methods, or approaches like data selection or data augmentation. To overcome these difficulties, we conceptualize a post-hoc bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify the class accuracy imbalance based on inference-time over-predictions and under-predictions.
To reduce COBias and rectify LLM outputs, we directly operate on the class probabilities obtained from in-context learning (ICL). We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to adjust class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights class probability with an optimal set of weights selected based on both metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP. Our key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the mathematical model.
  • Experiments show that DNIP achieves the best of both worlds of lower COBias and higher overall accuracy. For example, on Llama-2-13B, DNIP significantly reduces COBias (avg. $40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg. $59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg. $39.2\%\rightarrow 10.3\%$).
@@ 2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
2.1 Over-prediction and Under-prediction in Classification
[…]
In the era of Large Language Models (LLMs), over-prediction and under-prediction still persist in LLMs’ predictions, though one may not simply attribute such issues to the black-box training data or the model itself. The first obstacle is that the pretraining data’s source distribution for a subgroup of data may not be accessible, making it difficult to directly minimize differences between the source distribution and a downstream target distribution. What we normally have is the pretrained model’s output distribution for downstream classes. Secondly, the process of prompting an LLM, especially when the prompt is provided with in-context learned demonstrative examples, also brings uncertainty to predictions. Ablating the potential biases in prompting is non-trivial. More inconveniently, given the relatively large model sizes, it is computationally intensive, expensive, or even infeasible to tweak the LLM training pipeline or conduct LLM fine-tuning to fit debiasing training objectives.
To see how over-predictions and under-predictions hurt LLM prediction capabilities, we show in Table 2 an example of observed over-predictions and under-predictions in Llama-2-13B. We prompt the LLM on the AGNews news classification test set in a one-shot manner, with a single demonstrative example randomly selected from the training set. Llama-2-13B greatly over-predicts Business while under-predicting Technology, lowering the LLM’s capability for accurately predicting the Tech class (Tech’s accuracy: 19%). In general, there are 2129 predictions in Business but the number of ground-truth instances in Business is 1204; there are only 246 correct Tech predictions but 1240 ground-truth Tech instances. Over-/Under-prediction happens in more downstream tasks and larger LLMs (e.g., ChatGPT), as additionally show by examples in Appendix A. These surface observations tell us how an LLM could fail in predicting certain classes, suggesting that there is still space for improvements for LLMs to predict every class with equally good accuracy. The fact that this problem is difficult to completely eliminate from the source of data, together with the inconveniences in training/fine-tuning debiased LLMs, call for a more actionable inference-time bias metric quantifying the over-/under-predictions and an accompanying debiasing method.
@@ 2.1 Over-prediction and Under-prediction in Classification
Table 2: Significantly reduced over-predictions in class Business and improved test accuracy of the under-predicted class Tech for AGNews after DNIP.
2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
@@
In Table 2, we observe that Llama-2-13B greatly over-predicts Business while under-predicting Technology (Tech). Underneath the surface over-/under-prediction issue, we can see the most significant contributor is the 822 ground-truth instances of Technology, which are wrongly predicted as Business, leading to the model’s low accuracy (19%) in Technology [^1] It simultaneously leads to a fairly low precision score of 55% in Business. In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other. In other words, class accuracies are essential indicators for class performances. All current evaluation metrics are computed by Python scikit-learn. Note that by scikit-learn, class recall is the same as class accuracy.. On that note, we are able to precisely describe the surface over-prediction and under-prediction issue, by decoupling the problem into which output class that a ground-truth class is most wrongly predicted as, suggesting the model’s inference bias towards an output class given a ground-truth class. In the AGNews example, Business is what ground-truth Technology is most wrongly predicted as, suggesting the model’s inference bias to predict Technology as Business.
To quantify this inference bias, we first conceptualize odd classes. An odd class is an output class at inference time which holds the most wrong predictions of a ground-truth class $c_{i}$. That is, an odd class is always relative to a ground-truth class. For $N$ classes and $M$ test instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, and let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=c_{j}\},$ (1)
$odd_class|(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
[…]
@@
For this goal, it could be direct to use the rate of wrong predictions in an odd class as a measurement of its “hurt” to the accuracy of a ground-truth class, but minimizing these rates alone may not adequately mitigating class accuracy imbalances, which is an ultimate goal of reducing over-/under-predictions (More analyses on this measurement are in Appendix B.). A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class. For example, knowing that Business is odd to Tech, the 78% accuracy difference between Business (97%) and Tech (19%) can serve as a measurement of how much Tech is under-predicted. The 19% accuracy of Tech also implicitly reflects the many predictions in its odd class Business. It should be noted that measuring these accuracy differences sometimes fails to capture the hidden ground-truth-and-odd relationships. For instance, World is the odd class to ground-truth Business, but the -12% accuracy difference can not utter this fact. Despite this drawback, directly minimizing the accuracy differences will in return significantly reduce predictions in odd classes and improve under-predicted classes’ accuracies.
The Contextual Oddity Bias. To this end, we formulate the definition of Contextual Oddity Bias (COBias). First, the absolute value of the accuracy difference between a ground-truth class and its odd class is termed as COBiassingle. COBiassingle measures a model’s tendency to over-predict an answer that is caught as odd than the correct answer (or equivalently, a tendency to under-predict the correct answer). Notably, for LLMs, an odd answer may be a manifest of surface patterns from prompts or biases in the pretraining data. In addition, the odd classes in a same downstream task can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
@@
Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$. By few-shot prompting, the model’s output token probability over $N$ classes $c_{1},\dots,c_{N}$ is $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$. $\boldsymbol{p}_{m}$ is normalized over $N$ classes, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (3)
@@
where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances. Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances.
Moreover, besides the odd class, which is the most-frequent wrongly predicted class, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class. For a ground-truth class, there are the odd, the second odd, to the least odd classes. For example, for ground-truth Tech (1240 instances), Business is the odd (822), while World is the second odd (156), and Sports is the least odd (27). Then, we extend the accuracy difference between the single odd class and a ground-truth class to every pair of classes, and absolute value can be taken. This arrives at the ultimate the COBias metric as follows.
@@
This COBias metric directly reflects the inter-class imbalance in accuracy. Indeed, the absolute value of accuracy differences of some pairs in Equation 5 may mask the most frequent wrongly predicted class (the odd class). Nonetheless, it reflects a well-rounded measurement of all pairs of classes, whether involving odd classes or not, so it is a powerful metric when we strongly demand minimal class accuracy imbalance.
In a nutshell, the debiasing problem is minimizing COBias (Eq. 5) while maintaining or increasing the overall inference accuracy. This post-hoc bias evaluation metric will enable easy-to-operate inference-time/post-hoc output likelihood optimization, lifting the heavy computations required for training-wise data selection or model parameter updates.
@@ 3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, class probabilities, by directly minimizing the COBias metric while maximizing two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in during test time. An overview of the proposed DNIP method is shown in Figure 1.
@@ 3.1 The DNIP Objective
To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it multiplies by a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale. To find the optimal combination, the brute force strategy is an enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not run in polynomial time, calling for smarter methods.
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. Without loss of generality, we boil down the problem to selecting optimal weight indices towards minimizing the pairwise class accuracy imbalance, while maximizing the overall accuracy and the pointwise mutual information between class labels and predictions. Formally, we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
$\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (6)
The DNIP objective function and constraints are formally given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (7)
@@
$A_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\},\qquad i=1,\dots,N$ (9)
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},\qquad j=1,\dots,N$ (10)
@@
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively. In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.
The debiased class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi_{1}}p_{m1},\dots,\omega_{\xi_{N}}p_{mN})$ (12)
$y^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (13)
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class. The PMI term can also be seen as a constraint to penalize each individual class.

### v5

# COBias and Debias: Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

@@ Abstract
Large language models (LLMs) are good knowledge bases but struggle to perform equally well for all classes in text classification tasks. This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies. We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others. This class accuracy imbalance is difficult to solve from the root via better pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time combinatorial optimization. To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update. Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve it. Extensive evaluations on three LLMs across seven NLP classification tasks in different prompting settings show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
@@ 1 Introduction
The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience and reduce safety risks.
It could be difficult to disentangle all the factors that cause class accuracy imbalance. Resolving the issue from the root through pre-training or fine-tuning, leveraging data selection or data augmentation, is neither easy nor cost-effective. To overcome the issue, we conceptualize a bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify class accuracy imbalance from inference-time over-predictions and under-predictions. We aim to reduce COBias by directly rectifying the in-context learned LLM output class probabilities. We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to search for correction weights leveraging combinatorial optimization, without updating LLM parameters (Section 3). These weights adjust class probabilities towards lower COBias and higher overall accuracy. On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP.
Our key messages are as follows:
  • We introduce the COBias evaluation metric to quantify the class accuracy imbalance in language model predictions, where LLMs of different scales and families consistently exhibit a large COBias score.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) method to find class correction weights to correct ICL output class probabilities, which minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the nonlinear integer programming model.
  • DNIP achieves the best of both worlds of lower COBias and higher overall accuracy. For example, on Llama-2-13B, DNIP significantly reduces COBias (avg. $40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg. $59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg. $39.2\%\rightarrow 10.3\%$).
@@ 2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
Table 2: Significantly improved test accuracy of the under-predicted class Tech for AGNews after DNIP correction.
Over-predictions and under-predictions can hurt LLM prediction capabilities. To see how, Table 2 presents an example of observed over-predictions and under-predictions in Llama-2-13B. We prompt the LLM on the AGNews news classification [2], and find that Llama-2-13B greatly over-predicts Business while under-predicting Tech, leading to low accuracy in Tech (19%). Over-/Under-prediction happens in many downstream tasks and also in larger models such as ChatGPT (Appendix A), suggesting room for improvements so that LLMs can predict all classes with equally good accuracy.
@@
Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias. To introduce the Contextual Oddity Bias, we first elaborate the over-/under-prediction observations. In Table 2, the most significant contributor to the under-prediction in Tech (accuracy: 19%) is the 822 ground-truth Tech instances which are wrongly predicted as Business [^1] It simultaneously leads to a fairly low precision score of 55% in Business. In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other. In other words, class accuracies are essential indicators for class performances. All current evaluation metrics are computed by Python scikit-learn. Note that by scikit-learn, class recall is the same as class accuracy.. That is, the output class Business which a ground-truth class Tech is most wrongly predicted as can reflect the model’s inference bias towards an output class for the given ground-truth class. To precisely quantify this inference bias, we first conceptualize odd classes. An odd class is an output class which holds the most wrong predictions of a ground-truth class. That is, there is always an odd class to a ground-truth class. For $N$ classes and $M$ instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=i,\hat{y}_{m}=j\},$ (1)
$\textrm{odd}_{i}=\textrm{odd\_class}(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
@@
where $m=1,...,M,i,j=1,...,N,j\neq i$, $\mathbbm{1}\{\cdot\}=1$ if the condition holds and 0 otherwise. For example, odd_class(Tech) = Business. Then, we propose to quantify how much an odd class contributes to under-predicting a ground-truth class. One may immediately use the ratio of the number of odd class predictions over that of the ground-truth instances, e.g., use 822/1240 to measure the odd Business’s “hurt” to the ground-truth Tech, but minimizing these rates alone may not adequately mitigating class accuracy imbalances (More analyses on this measurement are in Appendix B.). A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class. For example, the 78% difference between Business (97%) and Tech (19%) quantifies how much Tech is under-predicted. The 19% also implicitly reflects the 822 wrong predictions in Business. It is noted that these accuracy differences sometimes fail to capture the ground-truth-and-odd relationships. For instance, World is the odd class to Business, but the former is 12% (97%-85%) below the latter. Despite the drawback, minimizing these differences will in return significantly reduce odd class predictions and improve under-predicted classes’ accuracies.
@@
The Contextual Oddity Bias. To this end, we propose Contextual Oddity Bias (COBias). First, we propose COBiassingle, the absolute value of the accuracy difference between a ground-truth class and its odd class. Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$. After obtaining few-shot prompted softmax probabilities over the vocabulary at the output posistion, we take the probabilities for the class tokens and normalize them to form class probabilities, i.e., $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$ is the normalized class probabilities for classes $c_{1},\dots,c_{N}$. Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
@@
where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances. Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances. COBiassingle measures a model’s tendency to under-predict the true answer and mostly mistake it as an odd answer. The odd classes in a same downstream task can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
Beyond the single odd classes, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class, i.e., the odd, the second odd, to the least odd classes. For example, for ground-truth Tech, Business is the odd, while World is the second odd, and Sports is the least odd. We extend $\textrm{COBias}_{\textrm{single}}$ to every pair of classes. This arrives at the ultimate the COBias metric as follows.
@@
COBias directly reflects the inter-class imbalance in accuracy. Although the absolute value of accuracy differences of some pairs in Equation 4 may mask the odd class, it reflects a well-rounded measurement and a powerful metric when minimal class accuracy imbalance is demanded.
@@ 3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Essentially, we aim to rectify biased few-shot prompted LLM class probabilities with correction weights, so that corrected class probabilities should lead to less class accuracy imbalance. With this goal, we propose an inference-time nonlinear integer programming (NIP) selection model to search for optimal correction weights from a weight scale that directly minimizes COBias while maximizing overall accuracy. Therefore, the correction weights are optimized by the NIP model on a set of given LLM output class probabilities with ground-truth labels. For a test instance, after obtaining its class probabilities, the learned correction weights are applied to correct them. An overview of the proposed Debiasing as Nonlinear Integer Programming (DNIP) method is shown in Figure 1.
@@ 3.1 The DNIP Objective
We first introduce a discrete $K$-point correction weight scale $\boldsymbol{\omega}$ to select correction weights from. We denote $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$, and $\omega_{k}=k/K$ is a value from the scale. For example, a $10$-point scale renders $\boldsymbol{\omega}=(0.1,\dots,1.0)$. To debias the $N$-dimensional $\boldsymbol{p}_{m}$, for each class $j$, $\boldsymbol{p}_{mj}$ multiplies a correction weight selected from $\boldsymbol{\omega}$ to obtain a corrected class probability.
Therefore, finding optimal correction weights for all $N$ classes becomes a combinatorial optimization problem: each class selects a weight from the $K$-point scale, so we need to find the optimal combination of the selections. To this end, we formulate the following correction weight selection model towards minimizing pairwise class accuracy differences and maximizing overall accuracy. Formally, we define integer selection variables $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely, $\xi_{n}=k$ if the $n$-th class selects $\omega_{k}$. As a result, the DNIP model can be given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+$
$\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (5)
@@
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},$
$j=1,\dots,N$ (7)
@@
where $\beta$ and $\tau$ are parameters that control the bias and the PMI terms respectively. The PMI term is introduced to maximize the pointwise mutual information between ground-truth class $j$ instances and predicted class $j$, which can be seen as a confidence level, or a constraint to penalize each individual class. $\hat{S}^{j}$ and $S^{j}$ denote the instances of prediction $j$ and true label $j$ respectively, $f(\hat{S}^{j})$ is the ratio between the number of instances with prediction $j$ and the total number of instances, similarly, $f({S}^{j})$ is the ratio between the number of instances labeled $j$ and the total number of instances, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of instances. In actual computations, we apply add-$\mu$ smoothing to the respective ratios in the numerator and denominator to smooth out zeros. The smoothing coefficient $\mu$ is selected along other parameters on a development set.
Intuitively, the first term directly minimizes error rates, the second term COBias directly minimize the inter-class accuracy differences, and the PMI term, is employed to further encourage predictions within a class to be close to the actual class. The corrected class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi^{\ast}_{1}}p_{m1},\dots,\omega_{\xi^{\ast}_{N}}p_{mN})$ (9)
$\hat{y}^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (10)